OpenStudy (anonymous):

Find the derivative of the function. y = e^x − e^(−x) / 5 y'=

OpenStudy (unklerhaukus):

\[y=e^x-{e^{-x}\over 5}\]\[y'=e^x+{e^{-x}\over5}\]

OpenStudy (anonymous):

\[y=\]\[(e ^{x}-e ^{-x})\div5\]

OpenStudy (anonymous):

i think my parentheses were confusing

OpenStudy (anonymous):

my apologies

OpenStudy (unklerhaukus):

oh i see now

OpenStudy (unklerhaukus):

\[y(x)={e^x-e^{-x} \over 5}\] \[y'(x)={e^x +e^{-x} \over 5}\]

OpenStudy (anonymous):

why did only the sign change?

OpenStudy (unklerhaukus):

well the thing about derivatives of e is they are easy \[(e^x)'=e^x\] \[(e^{ax})'=ae^{ax}\]

OpenStudy (unklerhaukus):

the second in exponential term a=-1